If the basis of a third order De are: y = 1,7=e- *.y=e6x. Then the corresponding De is: Select one: a. xy - 4x?y-5xy+y=0 b
Answer #1

Given that the basis anders of third DE que : ) Y = 1 → Y2 = x 1 Y3 = 25% We have we andea DE :) to find the correspo espondi

or y= (0-60** + 56ge5e) [:$(20x) = 2847] ly= -6e-² 563 e5x -C have Differentiating y with respect to x (-624-+563 25x) Ce-

y = 4, +(22-7 + (3 e 5% is solution of the DE (2) basis whose are Vi=) Hence, the corresponding DE Y2 = e-> and is Y3 = e52 -